# Stability of Continuous Time Quantum Walks in Complex Networks

### v2

@@ VI Conclusion
In this work, we investigate the stability of quantum systems by analyzing the dynamics of CTQWs on various network topologies under different decoherence. We define stability in terms of the preservation of quantum properties over time and examine how this preservation is influenced by the interplay between network structure and decoherence model. To analyze this, we consider both simple network structures, including star, cycle, and complete graphs, and complex topologies such as Erdős–Rényi, small-world, and scale-free networks.
@@
We evaluate the dynamics of CTQWs on these networks to examine the persistence of quantum properties, both in the absence of noise and under three decoherence models namely, Haken-Strobl, intrinsic decoherence and QSW. This systematic analysis allows us to explore how network structure and decoherence jointly influence CTQW stability. To quantify the stability, first we compute the node probability distribution of the quantum walker and the $\ell_{1}$-norm of coherence. Our simulations for node probabilities and $\ell_{1}$-norm of coherence confirm that under Haken-Strobl noise and QSW, being unital and Davies irreducible on connected networks (Appendix A), relax to the maximally mixed steady state. This contrasts with the intrinsic decoherence model, which allows for non-trivial steady states and localization. In addition to these primary measures, we also use fidelity, quantum–classical distance, and von Neumann entropy measures for providing a comprehensive characterization of the preservation of quantum properties.
@@
Based on these metrics, we find that stability varies significantly across decoherence models, intrinsic decoherence preserves coherence most effectively, followed by Haken–Strobl noise, while QSW result in the most rapid coherence loss. Our baseline scaling analysis on the star graph with initialization on center node showed that the influence of network size ($N$) is also highly decoherence model dependent: larger networks proved more coherence preservation under Haken-Strobl and intrinsic decoherence, but less stable under QSW. As expected, preservation of coherence uniformly decreased with increasing decoherence rates ($\gamma$ and $p$). Our results reveal that coherence preservation depends not only on the type of decoherence but also on the underlying topology and the specific node at which the quantum walker is initialized.
Delving deeper, we observe that under Haken-Strobl noise, cycle exhibit high coherence in the absence of noise but are highly vulnerable under decoherence. In contrast, star network and complete network demonstrate greater robustness under Haken–Strobl and intrinsic decoherence, as quantified by the $\ell_{1}$-norm of coherence. In the case of QSW, star network initialized at peripheral nodes retain coherence longer and delay the onset of uniform probability distribution compared to central hub initialization. Interestingly, complete and star networks initialized at central nodes exhibit identical probability distributions and identical stability measures under Haken-Strobl and intrinsic decoherence, reflecting the impact of local connectivity due to the initialization in the fully connected hub node.
For complex networks, when coherence preservation is quantified using $\ell_{1}$-norm of coherence indicates that scale-free networks are the most stable under Haken–Strobl but exhibit faster coherence decay than Erdős–Rényi networks under QSW dynamics. When stability is assessed using fidelity and von Neumann entropy, we observe distinct temporal behaviors across Erdős–Rényi, small-world, and scale-free networks. These results indicate that the presence of hubs in scale-free networks contributes to their enhanced stability. However, when using the quantum–classical distance as a stability measure, particularly under noiseless and intrinsic decoherence case, all network types yield nearly identical trajectories, making it difficult to distinguish relative stability. The interpretation of stability is inherently application dependent. For example, high entropy may impair the efficiency of quantum search algorithms [bose2000communication], while a large quantum–classical distance, indicating strong deviation from classical dynamics can benefit the design of optimized chiral quantum walks [frigerio2022quantum].
In general, our results reveal that stability is a non-trivial interplay between the network topology and the physical mechanism of the noise. High-connectivity features, such as the hubs in heterogeneous networks, provide stability against node based (Haken-Strobl) decoherence by inducing localization. However, these same features are a critical vulnerability to edge based (QSW) decoherence, where they accelerate decay of quantum coherence. Moreover, the value of coherence in noiseless scenario in such highly connected networks, whose initialization at high-degree nodes tends to promote localization but reduces coherence in noiseless conditions. This observation highlights a fundamental trade-off between localization and coherence. The implications of this trade-off are particularly relevant for quantum information processing: coherence can serve as a resource for quantum search algorithms based on quantum walks [shi2017coherence, su2018coherence], whereas high fidelity often indicative of localization can be exploited for robust quantum memories [chandrashekar2015localized]. Based on our analysis, there is significant scope for future research on stability in noisy environments by investigating additional structural properties of networks and considering the dynamics of multiple continuous-time quantum walkers (multi-CTQWs).
@@ Acknowledgments
ALJ acknowledges the University Grants Commission (UGC), India, for financial support. CM acknowledges support from the Anusandhan National Research Foundation (ANRF) India (Grants Numbers SRG/2023/001846 and EEQ/2023/001080). We are grateful to Dr. Luca Razzoli for helpful correspondence and for pointing us to his work [razzoli2022universality], which provided the analytical explanation for the universality we observed in the star initial localization on center node and complete network dynamics.
Appendix A Proof for Steady state
In this appendix we prove that the steady state of both the Haken-Strobl noise and QSW dynamics must relax to the unique maximally mixed state ($\rho_{ss}=\frac{\mathbb{I}}{N}$). This is demonstrated by verifying that their respective Liouvillians are unital and that the systems satisfy the condition for Davies irreducibility [davies1976quantum].
A.1 Steady state for Haken-Strobl Noise
@@
To guarantee that this steady state is unique, we invoke the concept of Davies irreducibility [davies1976quantum]. We utilize the criteria proposed by Zhang and Barthel [zhang2024criteria], which state that a system is Davies reducible if and only if there exists a non-trivial projection $\mathcal{P}$ where (where ${\mathcal{P}}\neq 0$ and $\mathcal{P}\neq\mathbb{I}$) that satisfies two conditions: (i) $(\mathbb{I}-\mathcal{P}){P}_{k}\mathcal{P}=0$ for all Lindblad operators ${P}_{k}$. (ii) $(\mathbb{I}-\mathcal{P})(i{H}+\frac{1}{2}\sum_{k}{P}_{k}^{\dagger}{P}_{k})\mathcal{P}=0$. If we can prove that the only projections $\mathcal{P}$ satisfying these conditions are $\mathcal{P}=0$ and $\mathcal{P}=\mathbb{I}$, the quantum system is irreducible. For the first condition:
@@
Since the eigen values of projector can only be 0 or 1, the only possible values are $c=0$ or $c=1$. Thus $\mathcal{P}=0$ or $\mathcal{P}=\mathbb{I}$ are the only possible projections. This satisfies the condition for irreducibility. Therefore, the Haken-Strobl Lindblad operators and irreducible and the system has a unique steady state, $\rho_{ss}=\frac{\mathbb{I}}{N}$ as shown in Eq. 15. This analytical result for steady state aligns with our simulations irrespective of network topology as shown in Sec. V.1.
A.2 Steady state for Quantum Stochastic Walk
@@
$\begin{split}\frac{d\rho(t)}{dt}=\mathcal{L}(\rho)=-(1-p)i[H,\rho(t)]+p\sum_{kj}\left(L_{kj}|k\rangle\langle j|\rho(t)L_{kj}|j\rangle\langle k|-\frac{1}{2}\left\{L_{kj}|j\rangle\langle k|L_{kj}|k\rangle\langle j|\rho(t)\right\}\right).\end{split}$ (23)
Checking whether $c\mathbb{I}$ can be a valid steady state for QSW as done for Haken-Strobl noise A.1:
$\begin{split}\mathcal{L}(c\mathbb{I})&=-(1-p)i[H,c\mathbb{I}]+p\sum_{kj}\left(L^{2}_{kj}|k\rangle\langle j|\,c\mathbb{I}\,|j\rangle\langle k|-\frac{1}{2}\left\{L^{2}_{kj}|j\rangle\langle k|k\rangle\langle j|,\,c\mathbb{I}\right\}\right)\\ &=p\sum_{kj}\left(cL^{2}_{kj}|k\rangle\langle k|-\frac{1}{2}\left\{L^{2}_{kj}|j\rangle\langle j|,\,c\mathbb{I}\right\}\right).\end{split}$ (24)
where, $\frac{1}{2}\left\{L^{2}_{kj}|j\rangle\langle j|,c\mathbb{I}\right\}=cL^{2}_{kj}|j\rangle\langle j|$Substitute this in Eq. 17 gives:
@@
As we did in  A.1, we use the Zhang and Barthel criteria to check for irreducibility. Applying the first condition $(\mathbb{I}-\mathcal{P}){P}_{kj}\mathcal{P}=0$:
@@
This is identical to Eq. 16 for the Haken-Strobl case. It immediately implies that for QSW, the projection operator $\mathcal{P}$ must be diagonal. Now consider the off-diagonal jump operators $k\neq j$. Taking the $il^{th}$ element of this:
@@
Since the graph is connected, either all diagonal elements are 1 ($\mathcal{P}=\mathbb{I}$) or all are 0 ($\mathcal{P}=0$). This proves that the QSW dynamics is Davies irreducible and has a unique steady state as in Eq. (26), $\rho^{ss}=\frac{\mathbb{I}}{N}$. This analytical result perfectly aligns with our numerical simulations in Sec. V.1 and  V.2, where we consistently observed convergence to this uniform distribution of probabilities with coherence decays to zero regardless of the network topology.
Appendix B Extended Stability Metrics for Simple Topologies
To obtain a more comprehensive picture of stability beyond the $\ell_{1}$-norm of coherence in Sec. V.1, we also analyze the quantum-classical distance, fidelity with the initial state, and von Neumann entropy.
@@ Appendix B Extended Stability Metrics for Simple Topologies
Indeed, these additional metrics confirm that the complete graph and the star network (hub-initialized) consistently demonstrate the highest stability, a finding that holds for both Haken-Strobl and intrinsic decoherence. This reinforces our earlier conclusion from the $\ell_{1}$-norm of coherence analysis, which had already identified these two same networks with high degree nodes as the most robust under the Haken-Strobl model. This is evidenced by their high quantum-classical distance and fidelity, along with the lowest von Neumann entropy, as shown in Fig. 9(a),(b),(c),(e),(f),(g) and Fig. 10(a),(b). In contrast, the cycle network tends to classicalize more rapidly, indicated by its low quantum-classical distance (Fig. 9(b),(c)). However, these metrics also reveal that for the QSW model, the stability rankings invert. The cycle network, previously the least stable, now displays a slower classicalization (a higher quantum calssical distance) than the highly-connected complete and star network with center node initialization (Fig. 9(d)). Under QSW, the star network with peripheral node initialization emerges as the most stable configuration, showing the slowest decay of quantum-classical distance, high fidelity, and low entropy (Fig. 9(d),(h) and Fig. 10(c)). This is consistent with our coherence analysis, which also identified this configuration as the most robust under QSW. This emphasizes the intricate interplay between network topology, the walker’s initial condition, and the specific mechanism of decoherence, as elaborated on in the main conclusion (Sec. VI).
@@
Figure 9: Quantum-classical distance and fidelity with the initial state for complete, cycle and star networks with size N = 10. Quantum-classical distance and fidelity with the initial state for noisless (a),(e) and three decoherence models: Haken-Strobl (b),(f), intrinsic Decoherence (c),(g) , and QSW (d),(h). The grey dashed line in (a) is theoretical value of quantum classical distance under no noise
@@
Figure 10: Time evolution of von Neumann entropy for Complete, Cycle and Star networks of size N = 10. Von Neumann entropy under Haken-Strobl (a), Intrinsic Decoherence (b), and QSW (c).
Appendix C Node Centrality Effects in Scale-Free Networks

### v3

@@ IV Conclusion and Discussion
In this work, we investigate the stability of CTQW by analyzing its dynamics on various network topologies under different decoherences. We define stability in terms of the preservation of quantum properties over time and examine how this preservation is influenced by the interplay between network structure and decoherence model.
@@
To quantify the stability, we first compute the node probability distribution of the CTQW and the $\ell_{1}$-norm of coherence. These measures confirm that, under Haken-Strobl and QSW decoherence, the system Liouvillians are unital and Davies-irreducible on connected networks ( B), and therefore relax to the maximally mixed steady state. This contrasts with the intrinsic decoherence model, which allows for non-trivial steady states and localization.
@@
Further, our results suggest that stability varies significantly across decoherence models, intrinsic decoherence preserves coherence most effectively, followed by Haken–Strobl noise, while QSW results in the most rapid coherence loss. Our results reveal that coherence preservation depends not only on the type of decoherence but also on the underlying topology and the specific node at which the quantum walker is initialized, specifically in heterogeneous networks.
Delving deeper, we find that sparse and relatively homogeneous networks, such as the cycle, Erdős-Rényi, and small-world topologies, exhibit stronger probability delocalization. In contrast, dense and heterogeneous networks, including the complete, star, and scale-free topologies, show pronounced localization at the initially occupied node in the noiseless and intrinsic decoherence cases, while simultaneously exhibiting higher coherence values.
Under Haken-Strobl noise, the hub-initialized star, scale-free, and complete networks retain coherence longer than the other topologies. However, in the QSW model, the peripheral-initialized star network, along with Erdős–Rényi and small-world networks, demonstrates comparatively longer coherence preservation. These trends are clearly reflected in both the probability distributions and the $\ell_{1}$-norm of coherence. Quantitative analysis based on fitting the decay of the $\ell_{1}$-norm of coherence further supports these observations. In particular, extracting the relaxation time from coherence decay provides a direct measure of decoherence loss and establishes a connection between stability timescales and underlying network structural properties.
In contrast, other stability quantifiers do not always provide a uniform characterization. For instance, fidelity and von Neumann entropy often indicate the highest stability for scale-free networks across all decoherence models. Similarly, the quantum-classical distance suggests enhanced stability of scale-free networks under Haken-Strobl and QSW dynamics. However, in the noiseless and intrinsic decoherence regimes, the quantum-classical distance yields nearly identical trajectories for all complex networks, making it difficult to clearly distinguish their relative stability.
Our results also support the previous studies that the interpretation of stability is inherently application dependent. For example, high entropy may impair the efficiency of quantum search algorithms [54], while a large quantum–classical distance, indicating strong deviation from classical dynamics, can benefit the design of optimized chiral quantum walks [55]. However, these same features are a critical vulnerability to edge-based QSW decoherence, where they accelerate the decay of quantum coherence.
Moreover, the value of coherence in a noiseless scenario, particularly in highly connected networks, whose initialization at high-degree nodes tends to promote localization but reduces coherence in the noiseless conditions. This observation highlights a fundamental trade-off between localization and coherence. This trade-off is particularly important for quantum information processing. In applications where coherence is the main resource, such as quantum search algorithms [56, 57], network topologies such as cycle, Erdős–Rényi, and small-world networks are more suitable because they maintain higher coherence. In contrast, for applications that rely on high fidelity and localization, such as robust quantum memories [58], star, complete, and scale-free networks are more appropriate.
The present study considers only Markovian single CTQW dynamics on static networks. However, the approach can be extended to include non-Markovian effects [59], time-dependent network structures [60, 61], and multiple interacting quantum walks (multi-CTQW) [62, 24, 63, 25]. These results can also be tested experimentally, particularly in photonic waveguide arrays and integrated photonic chips [64, 65, 63], where different network structures can be designed and observed.
@@ Acknowledgments
ALJ acknowledges the University Grants Commission (UGC), India, for financial support. CM acknowledges support from the Anusandhan National Research Foundation (ANRF) India (Grants Numbers SRG/2023/001846 and EEQ/2023/001080). We are grateful to Dr. Luca Razzoli for helpful correspondence and for pointing us to his work [48], which provided the analytical explanation for the universality we observed in the star initial localization on the center node and complete network dynamics.
Data availability
Data sharing is not applicable, as this study did not analyze any raw data.
Appendix A Influence of Network Size and Decoherence Rates in Simple Networks
Figure 8: Effect of network size and decoherence rate on the $\ell_{1}$-norm of coherence for a star graph with hub node initialization. Panels (a)–(c) show the coherence decay for different network sizes NN under (a) Haken–Strobl noise, (b) QSW, and (c) intrinsic decoherence, each evaluated at fixed decoherence rates ($\gamma=0.1$ or $p=0.1$). (d)–(f) The bottom row illustrates the coherence decay for various decoherence rates ($\gamma$ or $p$) for the three corresponding decoherence models on a network of fixed size $N=10$.
This appendix presents a scaling analysis of the star network with central node initialization, examining the dependence of $\ell_{1}$-norm of coherence on network size $N$ and decoherence rates $\gamma$ and $p$. As shown in Fig. 8(a), for Haken-Strobl noise the time taken for coherence decay increases with network size, showing that stability improves as $N$ increases. In contrast, under QSW dynamics, the opposite trend emerges. As observed in Fig. 8(b), the coherence decays more rapidly as $N$ increases, indicating that larger networks are slightly less stable under this decoherence mechanism. For intrinsic decoherence, Fig. 8(c) shows that as $N$ increases, the coherence saturates faster and to a higher steady-state coherence value, suggesting greater stability. Although smaller networks exhibit a slight peak before stabilizing, the reduced overall decay for larger $N$ shows that stability is enhanced. Next, we study the dependence on decoherence rates for a fixed network size of $N=10$. As shown in Fig. 8(d) for Haken-Strobl noise and Fig. 8(e) for QSW, stability uniformly decreases as the respective rates $\gamma$ and $p$ increase. For intrinsic decoherence Fig. 8(f) shows that while larger $\gamma$ value dampen initial oscillations and lead to faster saturation, the final steady-state coherence value is dependent on the network size, not on the decoherence rate.
Appendix B Proof for Steady state
In this appendix we prove that the steady state of both the Haken-Strobl noise and QSW dynamics must relax to the unique maximally mixed state ($\rho_{ss}=\frac{\mathbb{I}}{N}$). This is demonstrated by verifying that their respective Liouvillians are unital and that the systems satisfy the condition for Davies irreducibility [50].
B.1 Steady state for Haken-Strobl Noise
@@
To guarantee that this steady state is unique, we invoke the concept of Davies irreducibility [50]. We utilize the criteria proposed by Zhang and Barthel [51], which state that a system is Davies reducible if and only if there exists a non-trivial projection $\mathcal{P}$ where (where ${\mathcal{P}}\neq 0$ and $\mathcal{P}\neq\mathbb{I}$) that satisfies two conditions: (i) $(\mathbb{I}-\mathcal{P}){P}_{k}\mathcal{P}=0$ for all Lindblad operators ${P}_{k}$. (ii) $(\mathbb{I}-\mathcal{P})(i{H}+\frac{1}{2}\sum_{k}{P}_{k}^{\dagger}{P}_{k})\mathcal{P}=0$. If we can prove that the only projections $\mathcal{P}$ satisfying these conditions are $\mathcal{P}=0$ and $\mathcal{P}=\mathbb{I}$, the quantum system is irreducible. For the first condition:
@@
Since the eigen values of projector can only be 0 or 1, the only possible values are $c=0$ or $c=1$. Thus $\mathcal{P}=0$ or $\mathcal{P}=\mathbb{I}$ are the only possible projections. This satisfies the condition for irreducibility. Therefore, the Haken-Strobl Lindblad operators and irreducible and the system has a unique steady state, $\rho_{ss}=\frac{\mathbb{I}}{N}$ as shown in Eq. 15. This analytical result for steady state aligns with our simulations irrespective of network topology as shown in Sec. III.1.
B.2 Steady state for Quantum Stochastic Walk
@@
$\begin{split}\frac{d\rho(t)}{dt}=\mathcal{L}(\rho)=-(1-p)i[H,\rho(t)]+p\sum_{kj}\left(L_{kj}|k\rangle\langle j|\rho(t)L_{kj}|j\rangle\langle k|-\frac{1}{2}\left\{L_{kj}|j\rangle\langle k|L_{kj}|k\rangle\langle j|\rho(t)\right\}\right);\end{split}$ (23)
checking whether $c\mathbb{I}$ can be a valid steady state for QSW as done for Haken-Strobl noise B.1:
$\begin{split}\mathcal{L}(c\mathbb{I})&=-(1-p)i[H,c\mathbb{I}]+p\sum_{kj}\left(L^{2}_{kj}|k\rangle\langle j|\,c\mathbb{I}\,|j\rangle\langle k|-\frac{1}{2}\left\{L^{2}_{kj}|j\rangle\langle k|k\rangle\langle j|,\,c\mathbb{I}\right\}\right)\\ &=p\sum_{kj}\left(cL^{2}_{kj}|k\rangle\langle k|-\frac{1}{2}\left\{L^{2}_{kj}|j\rangle\langle j|,\,c\mathbb{I}\right\}\right);\end{split}$ (24)
where, $\frac{1}{2}\left\{L^{2}_{kj}|j\rangle\langle j|,c\mathbb{I}\right\}=cL^{2}_{kj}|j\rangle\langle j|$. Substitute this in Eq. 24 gives:
@@
As we did in  B.1, we use the Zhang and Barthel criteria to check for irreducibility. Applying the first condition $(\mathbb{I}-\mathcal{P}){P}_{kj}\mathcal{P}=0$:
@@
This is identical to Eq. 16 for the Haken-Strobl case. It immediately implies that for QSW, the projection operator $\mathcal{P}$ must be diagonal.
Now consider the off-diagonal jump operators $k\neq j$. Taking the $il^{th}$ element of this:
@@
Since the graph is connected, either all diagonal elements are 1 ($\mathcal{P}=\mathbb{I}$) or all are 0 ($\mathcal{P}=0$). This proves that the QSW dynamics is Davies irreducible and has a unique steady state as in Eq. (26), $\rho^{ss}=\frac{\mathbb{I}}{N}$. This analytical result perfectly aligns with our numerical simulations in Sec. III.1 and  III.2, where we consistently observed convergence to this uniform distribution of probabilities with coherence decays to zero regardless of the network topology.
Appendix C Linking Initialized Node’s Network Properties with Relaxation Time
Figure 9: Comparison of CTQW initialized node’s properties and relaxation time hierarchies. Panels (a)–(d) display specific properties of initialized node, which are compared against the characteristic relaxation times $\tau$ shown in panels (e)-(h). Specifically, panels (e) and (f) show the relaxation times for simple topologies (Table 1), while panels (g) and (h) correspond to the complex network ensembles (Table 2) under Haken-Strobl and QSW noise, respectively.
In this Appendix, we illustrate the relationship between properties of initialized node and the characteristic relaxation time $\tau$ (Eq. (12)), which is derived from the Kohlrausch parameters $\beta$ and $\lambda$ in Eq. (11).
For simple network topologies (cycle, star, and complete), the closeness centrality of the initialization node (Fig. 9(a)) follows the same hierarchy as the $\tau$ values observed under Haken-Strobl noise Fig. 9(e) and Table 1. In the QSW regime, however, the relaxation times in Fig. 9(f) align with the inverse density-weighted degree of the initial node $\frac{1}{k_{i}\rho}$ , as shown in Fig. 9(b). Here, $k_{i}$ represents the degree of the initial node and $\rho=\frac{2E}{N(N-1)}$ denotes the network density [66], where E and N are the total number of edges and nodes, respectively.
For complex network ensembles, these node properties are ensemble-averaged over the same realizations used to compute the relaxation times in Table. 2. We observe that the ensemble average of density-weighted degree of the initial node ($\langle k_{i}\rho\rangle$) as shown in Fig. 9(c) mirrors the hierarchy of $\langle\tau\rangle$ under Haken-Strobl noise (Fig. 9(g)). Conversely, for QSW decoherence, the average eccentricity of the initial node Fig. 9(d) aligns well with the behavior of $\langle\tau\rangle$ (Fig. 9(h)), further highlighting how specific initialized node properties capture the stability of the system under different decoherence mechanisms.
This hierarchy dictates how the properties of initially localized node affects the retention of coherence. Under Haken-Strobl noise, a node based decoherence mechanism, high closeness centrality and a high average density weighted degree shows high relaxation times, indicating that well-connected starting node can sustain coherence for longer. In contrast under QSW decoherence, which depends on both nodes and edges, the highest relaxation correlates with inverse of density weighted degree and eccentricity of the initial node. This suggests that nodes which posses fewer connectivity or topologically remote retain coherence for longer durations.
Appendix D Extended Stability Metrics for Simple Topologies
@@
To obtain a more comprehensive picture of stability beyond the $\ell_{1}$-norm of coherence in Sec. III.1, we also analyze the quantum-classical distance, fidelity with the initial state, and von Neumann entropy. Indeed, these additional metrics confirm that the complete graph and the star network (hub-initialized) consistently demonstrate the highest stability, a finding that holds for both Haken-Strobl and intrinsic decoherence. This reinforces our earlier conclusion from the $\ell_{1}$-norm of coherence analysis, which had already identified these two same networks with high degree nodes as the most robust under the Haken-Strobl model. This is evidenced by their high quantum-classical distance and fidelity, along with the lowest von Neumann entropy, as shown in Fig. 10(a),(b),(c),(e),(f),(g) and Fig. 11(a),(b). In contrast, the cycle network tends to classicalize more rapidly, indicated by its low quantum-classical distance (Fig. 10(b),(c)). However, these metrics also reveal that for the QSW model, the stability rankings invert. The cycle network, previously the least stable, now displays a slower classicalization (a higher quantum calssical distance) than the highly-connected complete and star network with center node initialization (Fig. 10(d)). Under QSW, the star network with peripheral node initialization emerges as the most stable configuration, showing the slowest decay of quantum-classical distance, high fidelity, and low entropy (Fig. 10(d),(h) and Fig. 11(c)). This is consistent with our coherence analysis, which also identified this configuration as the most robust under QSW. This emphasizes the intricate interplay between network topology, the walker’s initial condition, and the specific mechanism of decoherence, as elaborated on in the main conclusion (Sec. IV).
@@
Figure 10: Quantum-classical distance and fidelity with the initial state for complete, cycle and star networks with size N = 10. Quantum-classical distance and fidelity with the initial state for noisless (a),(e) and three decoherence models: Haken-Strobl (b),(f), intrinsic Decoherence (c),(g) , and QSW (d),(h). The grey dashed line in (a) is theoretical value of quantum classical distance under no noise.
@@
Figure 11: Time evolution of von Neumann entropy for Complete, Cycle and Star networks of size N = 10. Von Neumann entropy under Haken-Strobl (a), Intrinsic Decoherence (b), and QSW dechorence (c).
Appendix E Influence of Average Degree and Network Size in Complex Networks
In this appendix, we investigate the impact of the average degree $\langle k\rangle$ and network size $N$ on the temporal evolution of coherence across different network topologies.
To evaluate the effect of average degree, we examine the $\ell_{1}$-norm of coherence across a range of its values. As shown in Fig. 12(a)-(c), the peak magnitude of the $\ell_{1}$-norm of coherence under Haken-Strobl noise increases with $\langle k\rangle$ across Erdős–Rényi and small-world topologies, a trend that holds for most average degree values in scale-free networks as well. Notably, as $\langle k\rangle$ increases decay of coherence becomes more rapid for QSW decoherence for all network topologies as shown in Fig. 12(d)-(f). Furthermore, the intrinsic decoherence model depicted in Fig. 12(g)-(i) reveals a distinct qualitative shift in dynamics; while coherence saturates to a stable value for low connectivity ($\langle k\rangle\lessapprox 4$), systems with higher connectivity ($\langle k\rangle\gtrapprox 5$) exhibit a transient maximum followed by a decay toward a non-trivial constant steady state $\ell_{1}$-norm of coherence.
The influence of network size $N$ on complex topologies is presented in Fig. 13. Under Haken-Strobl noise (Fig. 13(a)-(c)), ncreasing $N$ leads to higher peak coherence and more rapid decay rates in Erdős–Rényi and small-world networks; however, scale-free networks exhibit relative invariance to system size, likely due to the dominant role of the central hub. For QSW decoherence (Fig. 13(d)-(f)), the coherence trajectories display a less discernible dependence on $N$, suggesting that coherence decay is primarily governed by local connectivity rather than global system size. In contrast, under intrinsic decoherence (Fig. 13(g)-(i)), a clear scaling trend emerges: the value at which coherence stabilizes, whether attained via direct saturation or following a transient peak, scales positively with $N$ for Erdős–Rényi and small-world topologies. An analogous trend observed in scale-free networks across some of the investigated system sizes ($N=50,60,70$).
Figure 12: Influence of average degree $\langle k\rangle$ on $\ell_{1}$-norm coherence in complex networks.Temporal evolution of coherence for Haken-Strobl noise in (a)-(c), QSW decoherence in (d)-(f), and intrinsic decoherence in (g)-(i) across Erdős–Rényi, small-world, and scale-free topologies ($N=100$). For the small-world (Watts–Strogatz) and scale-free (Barabási–Albert) models, only even values of $\langle k\rangle$ are considered to satisfy the generation constraints of these models.
Figure 13: Influence of network size $N$ on $\ell_{1}$-norm coherence in complex networks. The temporal evolution of coherence is presented for Haken-Strobl noise in (a)-(c), QSW decoherence in (d)-(f), and intrinsic decoherence in (g)-(i) across Erdős–Rényi, small-world, and scale-free topologies with average degree $\langle k\rangle=4$.
Appendix F Node Centrality Effects in Scale-Free Networks